OpenStudy (anonymous):

F(x)=x^2+6 and G(x)=x+8\x find (g o f) (-7)

hero (hero):

Problem: Given, f(x)= x^2+6 g(x) = (x+8)/x Find: g(f(-7)) Solution: In general, (g o f)(x) = g(f(x)) In this case, g(f(x)) = (x^2+6+8)/(x^2+6) = (x^2+14)/(x^2+6) Therefore, (g o f)(-7) = g(f(-7)) g(f(-7)) = ((-7)^2+14)/((-7)^2+6) = (49 + 14)/(49 + 6) = 63/55 Thus, g(f(-7) = 63/55